# In-2-4D: Inbetweening from Two Single-View Images to 4D Generation

### v2

@@ 5 Conclusion, Limitations, Future Work
Our method has some limitations. First, our method produces un-natural deformations when the in-between motion is extreme. Since the resulting videos are used to lift the object motion to the 3D space, the subtle movements may not look natural in 4D space. A promising direction for future work would be to extend this approach to incorporate specific motion trajectories or other 2D or 3D conditional signals in 4D motion generation to provide more realistic dynamism. Additionally, the 3D and 2D components do not currently interact in a way that allows mutual correction. Another valuable avenue for future research would be end-to-end training, enabling these two components to influence each other and produce more coherent results both in 2D and 3D.
Acknowledgements. This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant. We sincerely thank labmates from GrUVi and Taiya Lab for their valuable suggestions, numerous brainstorming sessions and proofreading of the draft.
Figure 9: Qualitative Results. Having only the first and last frames of a motion, we are able to generate moving 3D objects that can be seen at different views. Objects seen from different view directions are still plausible although no direct supervision signal is available.
Figure 10: Visual comparison between our method and the baselines. Our method produces less geometric and appearance artifacts in comparison with baselines II. More visual comparisons will be provided in the supplementary material.

### v3

@@ 5. Conclusion, Limitations, Future Work
Our method has some limitations. First, our method produces un-natural deformations when the in-between motion is extreme. Since the resulting videos are used to lift the object motion to the 3D space, the subtle movements may not look natural in 4D space. A promising direction for future work would be to extend this approach to incorporate specific motion trajectories or other 2D or 3D conditional signals in 4D motion generation to provide more realistic dynamism. Additionally, the 3D and 2D components do not currently interact in a way that allows mutual correction. Finally, since our pipeline relies on an image-to-3D module, we also inherit its limitations.If the module fails to accurately reconstruct an object, the resulting 4D output will exhibit corresponding artifacts.
Acknowledgement We thank all the anonymous reviewers for their insightful comments and constructive feedback. This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant.